Session: Feb 2002 (session 1)
Wavelength: 18. cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Jader Monari at Mon Feb 11 13:40:47 MET 2002) Success
Everthing OK!

Torun (Kaz Borkowski at Mon Feb 11 14:20:11 MET 2002) Minor failures
See note for this wavelength. No other known problems.

Jodrell_Bank (Alastair Gunn at Tue Feb 12 11:18:32 MET 2002) Minor failures
Serious problems with Lovell telescope azimuth drive. Therefore, recorded Mk2 telescope, but used schedule generated for the Lovell telescope. Azimuth drive problem eventually fixed. Schedule was stopped at 18:22 on day 41 (after DA193) to replace Lovell telescope in schedule. Started tape number 4 recording Lovell telescope.

Cambridge (Alastair Gunn at Tue Feb 12 11:19:08 MET 2002) Failure
Did not observe. Cambridge telescope link failed.

Effelsberg (A. Kraus at Wed Feb 13 11:41:03 MET 2002) Success

Westerbork (Tony Foley at Wed Feb 13 12:07:16 MET 2002) Minor failures
Large blocks of time out of hour angle range, so replaced with local calibrators. Local typo in editing this (by me) lost 07:18 to 07:57 on Monday Observing stopped at 09:58 UT monday as all scans after this point were out of hour angle range.

Westerbork (Tony Foley at Wed Feb 13 12:12:18 MET 2002) Minor failures
- supplemental (see Star Trek) several bursts of EMI at low end of the band (mostly BBC03 and BBC04)

Onsala (John Conway at Fri Feb 15 09:17:29 MET 2002) Minor failures
First 5 hours of expt lost due to high winds otherwise OK.